OpenStudy (anonymous):

Find the first 5 terms of the sequence.

OpenStudy (anonymous):

\[a _{1}=2, a _{n}=3a _{n-1}-1\] this is the sequence

OpenStudy (amistre64):

im sure if you follow the rule given it will be fine :)

OpenStudy (anonymous):

so does it start with 2?

OpenStudy (amistre64):

a1 = 2 a2 = 3(a1) - 1 yes

OpenStudy (anonymous):

Ok. thank you :) I just pretty much needed clarifications on this.

OpenStudy (amistre64):

a1 = 2 a2 = 3(2) - 1 = 5 a3 = 3(a2) - 1 3(5) - 1 = 14 etc ....

OpenStudy (amistre64):

yeah, understanding the notation itself is half the battle